OpenStudy (anonymous):

John is 250ft above ground on the top of a statue. He sees a ship in a harbor and measures the angle of depression as 18 degrees. Find the distance from the ship to the base of the statue.

OpenStudy (jonnyvonny):

To do this, we may make a triangle so we may visualize this  Now, we don't care about the hypotenuse, so don't worry about it. We have trig functions SOH-CAH-TOA. We may don't know the opposite length, but we do know the degree and the adjacent side, so tangent is the way to go. TOA means tan(theta) = opposite/adjacent so: \[\tan(\theta)=x/300\] We then may plug in values and solve for x. \[\theta= 18\rightarrow \tan(18) = x/300\rightarrow \tan(18)*300=x\] Plug that into your calculator (make sure it is in degree mode), then that's the answer. If you need more assistance with this, or anything else, feel free to ask. :)